A popular tent for packing purpose weight 3 pounds, 11 ounces. What is this mass in grams and in pictograms


1
Expert's answer
2021-04-28T06:17:24-0400

3 pounds =[ 3 × 16 ] = 48 ounces

 

Total = 48 + 11=59 ounces


In grams = 59 × 28.35 = 1672.65 grams


In pictograms = 1672.65/1000 = 1.67265